# Learning normalized image densities via dual score matching

### v1

@@ Appendix B Proofs and derivations
B.1 Time score matching (proof of eq. 4)
@@
Note that the derivation exactly matches that of the Miyasawa-Tweedie identity by replacing $\partial_{t}$ with $\nabla_{y}$, and does not make any assumptions about the form of $p(y|x,t)$. Restricting to additive Gaussian noise, $-\log p(y|x,t)=\frac{1}{2t}{\mathopen{}\mathclose{{}\left\lVert y-x}\right\rVert}^{2}+\frac{d}{2}\log(2\pi t)$, and substituting into eq. 25 gives the “time score-matching identity of eq. 4:
@@ B.2 Energy architecture
We remark that if $s_{\theta}$ is homogeneous, then $U_{\theta}$ is quadratically homogeneous ($U_{\theta}(\lambda y,t)=\lambda^{2}U_{\theta}(y,t)$). Note that this does not correspond to enforcing (asymmetric) Gaussian one-dimensional marginals $\langle y,u\rangle$. Rather, this enforces that the distribution of $\langle y,u\rangle$ conditioned on the orthogonal projection $y-\langle y,u\rangle u/{\mathopen{}\mathclose{{}\left\lVert u}\right\rVert}^{2}$ is (asymmetric) Gaussian.
B.3 Effective dimensionality (equivalence between eqs. 11 and 12)
@@
Combining eqs. 31 and 35 into eq. 30 and integrating by parts twice, we have
@@
where we have used Stein’s lemma in the last step. We finally combine eqs. 36 and 39 to obtain
@@
We define the common value in eq. 40 to be $d_{\rm eff}(x,t)$. It is related (but not equal) to dimensionality measures estimated from the singular values of the Jacobian of a denoiser (Mohan* et al., 2020; Horvat and Pfister, 2024). The limit of $d_{\rm eff}(x,t)$ when $t\to 0$ has appeared in the literature under the name of local intrinsic dimensionality using (approximate) likelihood (LIDL) (Tempczyk et al., 2022; Stanczuk et al., 2022; Kamkari et al., 2024). Its average over images $x$ is equal to the MMSE dimension of Wu and Verdú (2011).
@@ C.1 Energy network architecture
Our UNet architecture is composed of $3$ encoder blocks, a middle block, and $3$ decoder blocks. Each block is itself composed of $3$ layers, each a sequence of bias-free convolution, normalization, and non-linearity, for a total of $21$ layers. The first convolutional layer of each encoder block but the first and the middle block has a stride of $2$ (downsampling, in both vertical and horizontal directions), and the last convolutional layer of each decoder block (except for the last one and the middle block) is transposed with a stride of $2$ (upsampling). The output of each encoder block is concatenated to the input of the corresponding decoder block (which comes from the output of the corresponding encoder block, via a “skip connection”). The number of channels is doubled in each block, starting from a base value of $64$ at the coarsest scale. We replace ReLUs with GeLUs to ensure that $\nabla_{y}U_{\theta}$ is differentiable. Thus, $U_{\theta}$ is only approximately quadratically homogeneous.
@@ C.2 Training hyper-parameters
To summarize Section 2.2, the dual score matching training objective is
@@
[…]
@@ C.3 Additional details
Gaussian scale mixture example (Figure 1).
@@
where the functions $a_{i},b_{i}$ are computed by a $5$-layer MLP with a hidden dimension of $256$ that takes $\log(t+t_{\min})$ as input. This network is trained either with single (space) score matching or with dual score matching, both across noise levels $t\in[t_{\min},t_{\max}]$. Training is otherwise similar to the ImageNet64 experiments (see Section C.2), for $20,000$ training steps with a batch size of $512$ and an initial learning rate of $0.0001$, over noise levels from $t_{\min}={10}^{-2}$ to $t_{\max}={10}^{2}$. We note that the energy learned after single score matching training is stochastic: as the relative energy between the two mixture components is not constrained by the data, its value is determined by the random initialization, so that rerunning the experiment will lead to a different value.
Histogram of log probabilities (Figure 3).
The Gumbel fit is calculated by maximizing likelihood. The obtained parameters (in decibels/dimension) are $9.57$ for the location and $3.17$ for the scale. Equivalently, $p(x)$ follows a Fréchet distribution, with a scale parameter equal to ${9.05}^{d}$ and shape parameter equal to ${1.37}/d$, where $d=64\times 64=4096$ is the dimension of ImageNet64 grayscale images.
Computing dimensionality (Figure 5).
Equations 11 and 12 provide two ways to estimate effective dimensionality from a learned energy model. If the model is exact, $U_{\theta}(y,t)=U(y,t)$, then they coincide (more generally, they coincide when the model satisfies the diffusion equation), but in general they only approximately coincide. For numerical stability, we found it preferable to use the version defined from the mean squared error of the underlying denoiser (eq. 11). This guarantees non-negative dimensionalities, and also has the advantage of being an upper bound on the true dimensionality (as any denoiser yields an upper bound on the minimum MSE).
Appendix D ImageNet images according to their probability
@@
Figure 8: Images in ImageNet64 (test set) with linearly-spaced log probabilities.

### v3

@@ Appendix B Proofs and derivations
B.1 Time score matching (proof of eq.˜4)
@@
Note that the derivation exactly matches that of the Miyasawa-Tweedie identity by replacing $\partial_{t}$ with $\nabla_{y}$, and does not make any assumptions about the form of $p(y|x,t)$. Restricting to additive Gaussian noise, $\log p(y|x,t)=\frac{1}{2t}{\mathopen{}\mathclose{{\left\lVert y-x}}\right\rVert}^{2}+\frac{d}{2}\log(2\pi t)$, and substituting into eq.˜25 gives the “time score-matching” identity of eq.˜4:
@@ B.2 Energy architecture
We also remark that if $s_{\theta}$ is homogeneous, then $U_{\theta}$ is quadratically homogeneous ($U_{\theta}(\lambda y,t)=\lambda^{2}U_{\theta}(y,t)$). Note that this does not correspond to enforcing (asymmetric) Gaussian one-dimensional marginals $\langle y,u\rangle$. Rather, this enforces that the distribution of $\langle y,u\rangle$ conditioned on the orthogonal projection $y-\langle y,u\rangle u/{\mathopen{}\mathclose{{\left\lVert u}}\right\rVert}^{2}$ is (asymmetric) Gaussian.
B.3 Effective dimensionality (equivalence between eqs.˜11 and 12)
@@
Combining eqs.˜31 and 35 into eq.˜30 and integrating by parts twice, we have
@@
where we have used Stein’s lemma in the last step. We finally combine eqs.˜36 and 39 to obtain
@@
We define the common value in eq.˜40 to be $d_{\rm eff}(x,t)$. It is related (but not equal) to dimensionality measures estimated from the singular values of the Jacobian of a denoiser (Mohan* et al., 2020; Horvat and Pfister, 2024). The limit of $d_{\rm eff}(x,t)$ when $t\to 0$ has appeared in the literature under the name of local intrinsic dimensionality using (approximate) likelihood (LIDL) (Tempczyk et al., 2022; Stanczuk et al., 2022; Kamkari et al., 2024). Its average over images $x$ is equal to the MMSE dimension of Wu and Verdú (2011).
@@ C.1 Energy network architecture
Our UNet architecture $s_{\theta}$ is composed of $3$ encoder blocks, a middle block, and $3$ decoder blocks. Each block is itself composed of $3$ layers, each a sequence of bias-free convolution, normalization, and non-linearity, for a total of $21$ layers. The first convolutional layer of each encoder block but the first and the middle block has a stride of $2$ (downsampling, in both vertical and horizontal directions), and the last convolutional layer of each decoder block (except for the last one and the middle block) is transposed with a stride of $2$ (upsampling). The output of each encoder block is concatenated to the input of the corresponding decoder block (which comes from the output of the corresponding encoder block, via a “skip connection”). The number of channels is doubled in each block, starting from a base value of $64$ at the coarsest scale. We replace ReLUs with GeLUs to ensure that $\nabla_{y}U_{\theta}$ is differentiable. Thus, $U_{\theta}$ is only approximately quadratically homogeneous. As a result, and because training does not result in an exactly conservative $s_{\theta}$, $\nabla_{y}U_{\theta}(y,t)$ should be computed directly as opposed to using $s_{\theta}$.
@@ C.2 Training hyper-parameters
To summarize Section˜2.2, the dual score matching training objective is
@@
We use the ImageNet64 dataset (Chrabaszcz et al., 2017), with (only) horizontal flips as data augmentation. The models used in Tables˜1 and 2 and for the generalization experiment in Figure˜2 are trained on color images, while the model used in Figures˜3, 4 and 5 is trained on grayscale images. Pixel values are rescaled to $[0,1]$ by dividing by $255$. All models are trained for $1\text{M}$ steps, with a batch size of $128$. We use the Adam optimizer with default parameters and an initial learning rate of $0.0005$ (except for the generalization experiments which used a learning rate of $0.0002$) that is halved every $100,000$ steps. All models are trained on a single NVIDIA H100 GPU, which takes about $5$ days for ImageNet64.
@@ C.3 Additional details
Gaussian scale mixture example (Figure˜1).
@@
where the functions $a_{i},b_{i}$ are computed by a $5$-layer MLP with a hidden dimension of $256$ that takes $\log(t+t_{\min})$ as input. This network is trained either with single (space) score matching or with dual score matching, both across noise levels $t\in[t_{\min},t_{\max}]$. Training is otherwise similar to the ImageNet64 experiments (see Section˜C.2), for $20,000$ training steps with a batch size of $512$ and an initial learning rate of $0.0001$, over noise levels from $t_{\min}={10}^{-2}$ to $t_{\max}={10}^{2}$. We note that the energy learned after single score matching training is stochastic: as the relative energy between the two mixture components is not constrained by the data, its value is determined by the random initialization, so that rerunning the experiment will lead to a different value.
Histogram of log probabilities (Figure˜3).
The Gumbel fit is calculated by maximizing likelihood. The obtained parameters (in decibels/dimension) are $9.57$ for the location and $3.17$ for the scale. Equivalently, $p(x)$ follows a Fréchet distribution, with a scale parameter equal to ${9.05}^{d}$ and shape parameter equal to 1.37/d{1.37}/d, where $d=64\times 64=4096$ is the dimension of ImageNet64 grayscale images.
Computing dimensionality (Figure˜5).
Equations˜11 and 12 provide two ways to estimate effective dimensionality from a learned energy model. If the model is exact, $U_{\theta}(y,t)=U(y,t)$, then they coincide (more generally, they coincide when the model satisfies the diffusion equation), but in general they only approximately coincide. For numerical stability, we found it preferable to use the version defined from the mean squared error of the underlying denoiser (eq.˜11). This guarantees non-negative dimensionalities, and also has the advantage of being an upper bound on the true dimensionality (as any denoiser yields an upper bound on the minimum MSE).
Appendix D Additional results
D.1 ImageNet images according to their probability
@@
Figure 8: Example images from ImageNet64 (test set) with linearly-spaced log probabilities, ordered from low to high.
D.2 Distribution of log probabilities on other datasets
To evaluate the pervasiveness of the Gumbel-distributed log probability values, we trained energy models on CelebA (grayscale and color) and CIFAR-10, with an identical architecture, objective, and hyper-parameters as described in Section˜2 and Appendix˜C. Results are presented in Figures˜9, 10 and 11. In all cases, log-probability histograms share qualitative features with the Gumbel distribution, namely its skewness towards high-probability values, although the fits are less precise due to the presence of low- or high-probability outliers.
Figure 9: Histogram of log probabilities of test images in the CelebA dataset (grayscale). Color-coded arrows indicate values for the example images on the right. The maximum-likelihood fit by a Gumbel distribution (red line) is perturbed by the low-probability outliers, some of which correspond to corrupted images.
Figure 10: Histogram of log probabilities of test images in the CelebA dataset (color). Color-coded arrows indicate values for the example images on the right. The maximum-likelihood fit by a Gumbel distribution (red line) is perturbed by the low-probability outliers.
Figure 11: Histogram of log probabilities of test images in the CIFAR-10 dataset. Color-coded arrows indicate values for the example images on the right. The maximum-likelihood fit by a Gumbel distribution (red line) is perturbed by high-probability outliers, which correspond to grayscale images.
D.3 Log probabilities of interpolated images
Figure 12: Log probability of interpolated images. For each pair $(x_{i},x_{i+1})$ of consecutive example images (see Figure˜3), we consider a linear path between them, $\alpha\in[0,1]\mapsto(1-\alpha)x_{i}+\alpha x_{i+1}$ (shown in the right panel). In the left panel, we plot the log probability of the resulting images $\log p_{\theta}((1-\alpha)x_{i}+\alpha x_{i+1})$ as a function of the interpolated log probability $\alpha\log p_{\theta}(x_{i})+(1-\alpha)\log p_{\theta}(x_{i+1})$. For medium- and high-probability images, this leads to a sharp decrease in probability even for $\alpha$ close to 0 or 11: the superposition of two natural images is not a natural image. However, for low-probability images, this surprisingly leads to a slight increase in probability: the interpolation between two densely textured images is another densely textured image but with slightly lower contrast, leading to higher probability values (as shown in Figure˜4).
D.4 Optimizing log probability
Figure 13: For each initial image (middle row), we perform 1,0001,\!000 steps of gradient ascent (top row) or descent (middle row) on the log probability with a fixed step size of ${10}^{-2}/d$, clipping pixel intensity values to $[0,1]$ at each step. Due to numerical instabilities around high-probability images, we select the iterate with the highest or lowest probability. Resulting log probability values, and their difference with the starting value, are shown above or below the final images. Maximizing log probability leads to generally smoother images (although with some visual artifacts), except for sufficiently high-probability images, where gradient ascent fails to increase the probability. Minimizing log probability leads to noisier images, with regular geometric patterns for high-probability images.